OpenStudy (anonymous):

What is the vertex of the graph of y=-|2x+4|+3?

OpenStudy (anonymous):

you are looking for \((a,b)\) and the \(a\) is the number that will make \(2x+4=0\)

OpenStudy (anonymous):

the \(b\) is what get if \(|2x+4|=0\) namely \(3\)

OpenStudy (nathan917):

This term –|2x + 4| is always negative, therefore y= 3 is its maximum. y=3 when 2x+4=0, or x=-2 vertex(-2,3)